### v1

@@ 5 Experimental Results
To evaluate the performance of POCO, we compare it with state-of-the-art NCO methods and strong traditional solvers on typical COP benchmarks with various problem shapes and distributions. Performance evaluation is based on the gap metric $\frac{g(\bm{y})-g(\bm{y}^{*})}{g(\bm{y}^{*})}\times 100\%$ between the obtained solution $\bm{y}$ and known optimal solution $\bm{y}^{*}$, where a lower gap indicates better performance. The best results are highlighted in bold. We also report total solving time for each instance group. Experiments were conducted on a Linux system with an NVIDIA TITAN Xp GPU and an Intel(R) Xeon(R) E5-2680 CPU. Our code will be made publicly available.
5.1 Job-shop Scheduling Problem
Neural Model. Each JSP instance is represented as a disjunctive graph, a standard representation for scheduling problems. For details of disjunctive graphs, see Appendix A. We employ a neural model, named MGL, that combines a multi-layer graph attention network (GAT) Veličković et al. (2018) encoder for computing node embeddings with a long short-term memory (LSTM) based Hochreiter & Schmidhuber (1997) context-attention decoder for predicting action probabilities using both embedding and context features. The complete architecture is detailed in Appendix B.
Training & Test. For evaluation, we use three standard JSP benchmarks: Taillard’s (TA) Taillard (1993), Lawrence’s (LA) Lawrance (1984), and Demirkol’s (DMU) Demirkol et al. (1998). Each benchmark contains 8 different shapes with 10 instances per shape, except LA which has 5 instances per shape. We generate a training dataset of 30000 instances following SLIM Corsini et al. (2024), consisting of 6 shapes $(n\times m)$ in $\{10\times 10,15\times 10,15\times 15,20\times 10,20\times 15,20\times 20\}$ with 5000 instances per shape. During training, we generate additional 100 different instances per shape from the same shape set for validation. We employ the Adam optimizer Kingma & Ba (2014) with learning rate $\eta=0.0002$ and train the NCO model for 20 epochs. We set the solution number of hybrid rollout $B=256$, the number of filtered solutions $K=16$, batch size of $D=1$. During inference, we adopt both greedy rollout and sampling rollout with $B^{\prime}$ solutions.
[…]
@@ 5.1 Job-shop Scheduling Problem
Exact Solver RL-based Improvement Traditional PDR RL SLL POCO $B^{\prime}=128$ $B^{\prime}=512$
@@ 5.2 Traveling Salesman Problem
Neural Model. We adopt the same model as POMO Kwon et al. (2020), comprising an encoder with 6 Transformer layers and a decoder with a multi-head attention layer.
Training & Test. Following the NCO literature, we evaluate POCO on randomly generated instances with $n=20/50/100$ (denoted as TSP20/50/100), using models trained on corresponding problem shapes. Additionally, we assess POCO’s generalization capability on the out-of-distribution TSPLIB benchmark Reinelt (1991). We adopt the same hyperparameter configuration as POMO. During training, the number of filtered solutions is set to $K=8$. For TSP20 and TSP50, we set the hybrid rollout solution number $B=128$ and batch size $D=64$. For TSP100, we set $B=256,K=16$ and $D=48$ due to memory limit. Training instances are generated randomly. For inference, we adopt greedy rollout with multiple start nodes and $\times$8 instance augmentation (denoted as aug.), consistent with POMO.
Baselines. We assess POCO against both traditional methods and neural constructive methods. The traditional methods include Gurobi, an exact solver, and LKH3, a powerful problem-specific heuristic. The neural constructive methods include POMO Kwon et al. (2020), a widely-adopted RL-based backbone for advanced methods, and SLIM Corsini et al. (2024), a current state-of-the-art SLL-based method.
Results on TSP. As shown in Table 2, POCO outperforms other neural baselines while delivering competitive solutions against traditional solvers, despite the latter requiring substantially more computational time. Notably, POCO achieves superior performance over POMO on TSP100, even with fewer training epochs (700 vs. 2000). Generalization results on TSPLIB, presented in Table 3, demonstrate POCO’s significant advantages over other neural baselines when generalizing to out-of-distribution instances.
@@
POCO (aug.) 3.83 0.00 3.6s 5.69 0.01 6.6s 7.75 0.04 18.1s
@@ 5.3 Flexible Job-shop Scheduling Problem
Neural Model. For a FJSP instance represented as a disjunctive graph, we adopt the MGL model (see Appendix B) to compute action probabilities.
Training & Test. For evaluation, we use FJSP instances from the LA benchmark Lawrance (1984). The benchmark includes e-data, r-data, and v-data, where each operation can be allocated to 1-2 machines, 1-3 machines, and 1-$m$ machines, respectively. Following Song et al. (2023), we generate 25000 FJSP instances with 5 different shapes: {$10\times 5\times 5$, $10\times 10\times 10$, $15\times 10\times 10$, $20\times 5\times 5$, $20\times 10\times 7$}. For each shape, 2500 instances belong to r-data and 2500 to v-data. We set $B=128$, $K=16$, and $D=1$ during training, and employ both greedy rollout and sampling rollout with $B^{\prime}$ solutions during inference.
Baselines. For FJSP, we compare POCO with both representative traditional PDRs and state-of-the-art RL-based constructive methods: DNN Yuan et al. (2024) using the actor-critic framework, HG Song et al. (2023) utilizing heterogeneous graphs for instance representation, and RS Ho et al. (2024) employing residual scheduling to remove finished operations.
Results on FJSP Benchmarks. As shown in Table 4, with greedy rollout, POCO significantly outperforms most baselines, only marginally falling behind RS on LA e-data. When all methods use sampling rollout with 100 solutions, POCO even achieves the best performance across all cases. Notably, increasing the number of sampled solutions consistently reduces the optimality gap.
Figure 2: Training curves of different training paradigms for JSP.
@@
Traditional PDR RL POCO $B^{\prime}$=100 $B^{\prime}$=256 $B^{\prime}$=512
Benchmarks SPT MOR MWR DNN HG RS POCO HG RS POCO POCO POCO
@@
5.4 Ablation Study
[…]
Figure 3: Training curves of different training paradigms for TSP.
Effectiveness of the Preference Pair Construction Method. To validate our three-step preference pair construction method, we replace each step with a simpler alternative. We substitute sampling rollout for hybrid rollout, replace uniform filtering with random, top-$K$, or bottom-$K$ filtering, and use full permutation pairing instead of best-anchored pairing while maintaining the same total number of pairs. As shown in Table 5, POCO’s performance deteriorates without (w/o) any of these components, demonstrating the necessity of our design. Notably, top-$K$ filtering leads to significantly worse performance, highlighting the importance of filtering worse solutions to create sufficient preference differences in pair construction. Additional analyses of hybrid rollout effectiveness are provided in Appendix E.
@@ 5.4 Ablation Study
Shape Sampling Rollout Random Top-$K$ Bottom-$K$ Full Permutation Pairing POCO
@@
Superiority of the POCO Loss. To evaluate our POCO loss with a preference-based scaling factor, we compare it with other preference optimization losses: the classic DPO loss ($\mathcal{L}_{DPO}$) Rafailov et al. (2024), the popular SimPO loss ($\mathcal{L}_{SimPO}$) Meng et al. (2024), and a variant of POCO loss without the scaling factor ($\mathcal{L}_{POCO^{-}}$). Complete loss formulations are provided in Appendix C. For DPO, which requires a reference model to prevent excessive policy deviation, we use the old model from 10 episodes prior, similar to PPO, with the common hyperparameter setting $\beta=0.1$. For SimPO, we follow their standard parameters with $\beta=2$ and $\gamma=1$. As shown in Table 6, POCO achieves the best performance across all benchmarks for both $B^{\prime}=128$ and $B^{\prime}=512$, with particularly significant improvements on the DMU benchmark, demonstrating the effectiveness of our loss design and its preference-based scaling factor.
@@
Benchmark $\mathcal{L}_{SimPO}$ $\mathcal{L}_{DPO}$ $\mathcal{L}_{POCO^{-}}$ $\mathcal{L}_{POCO}$ $\mathcal{L}_{SimPO}$ $\mathcal{L}_{DPO}$ $\mathcal{L}_{POCO^{-}}$ $\mathcal{L}_{POCO}$
@@ 5.5 Hyperparameter Study
POCO has two crucial hyperparameters: the solution number of hybrid rollouts $B$ and the number of filtered solutions $K$. We analyze their individual effects by varying $B\in\{32,64,128,256,512\}$ and $K\in\{4,8,16,32\}$. Additionally, we explore their interaction by maintaining a fixed ratio $B/K=16$ while scaling both parameters $K\times B\in\{4\times 64,8\times 128,16\times 256,32\times 512\}$.
Effect of the Solution Number of Hybrid Rollouts. As shown in Figure 4(a), increasing the number of sampled solutions during training, i.e., larger $B$, improves solution quality by enhancing the probability of collecting higher-quality solutions. However, GPU memory consumption grows with $B$ due to parallel computation, raising costs. While the performance improves significantly up to $B=256$, further increases yield diminishing returns despite rising computational costs, making $B=256$ a reasonable choice.
@@ 6 Conclusion
In this work, we present POCO, a preference optimization-based training paradigm for NCO. By introducing tailored preference pair construction and a novel loss function for COPs, POCO achieves higher sample efficiency than mainstream RL and recent SLL paradigms. Extensive experiments on JSP, TSP, and FJSP demonstrate POCO’s superior performance over state-of-the-art neural constructive methods, while requiring significantly less time to deliver solutions competitive with traditional problem-specific iterative heuristics. POCO requires neither expensive labels nor specialized design of the Markov Decision Process, making it easy to use in practice. More importantly, it establishes a general training paradigm that can be readily applied to various NCO models for solving different COPs.
Although POCO demonstrates superior sample efficiency compared with RL and SLL, one limitation is that it still requires a relatively large number of rollout solutions, similar to SLIM. This incurs moderate costs in collecting high-quality solutions for effective model learning. In future work, we will explore efficient ways to obtain high-quality solutions that facilitate model training. One promising direction is to leverage problem invariance and solution symmetry to efficiently generate diverse high-quality training data. Another direction is to efficiently enhance solution quality by incorporating problem-specific heuristics during training, providing better learning signals for the model.
@@ A.1 Job-shop Scheduling Problems
Disjunctive Graph. JSP can be represented using a disjunctive graph $G=(V,A,E)$. In this graph, the node set $V=\{O_{j}\;|\;O_{j}\in\mathcal{O}\}$ represents operations, the directed edge set $A=\{O_{l_{ij}}\rightarrow O_{i(j+1)}\;|\;O_{l_{ij}},O_{i(j+1)}\in\mathcal{O}\}$ indicates precedence constraints between successive operations ($O_{l_{ij}}\rightarrow O_{l_{i(j+1)}}$), and the disjunctive (undirected) edge set $E$ connects operations performed on the same machine. A feasible solution is obtained by assigning directions to the undirected edges in $E=\{O_{j}\leftrightarrow O_{j^{\prime}}\;|\;O_{j},O_{j^{\prime}}\in\mathcal{O},M_{j}=M_{j^{\prime}}\}$, resulting in a directed acyclic graph.
Features of JSP. Following previous works Zhang et al. (2020); Corsini et al. (2024), we define two types of features. The first type consists of static state features $\bm{s}_{j}$ assigned to each operation $O_{j}$ in the node set $V$, while the second type comprises contextual features $\bm{c}_{i}$ based on job $J_{i}$ information at step $t$. The state features $\bm{s}_{j}$ are fed into the encoder to compute embeddings for each node (operation), while the contextual features $\bm{c}_{i}$ are combined with these embeddings to jointly contribute to the prediction. Table 7 details the state features $\bm{s}_{j}\in\mathbb{R}^{15}$, and Table 8 details the contextual features $\bm{c}_{i}\in\mathbb{R}^{11}$.

### v3

@@ 5 Experimental Results
To evaluate the performance of BOPO, we compare it with state-of-the-art NCO methods and strong traditional solvers on typical COP benchmarks with various problem shapes and distributions. Performance evaluation is based on the gap metric $\frac{g(\bm{y})-g(\bm{y}^{*})}{g(\bm{y}^{*})}\times 100\%$ between the obtained solution $\bm{y}$ and known optimal solution $\bm{y}^{*}$, where a lower gap indicates better performance. The best results are highlighted in bold. We also report total solving time for each instance group. Experiments were conducted on a Linux system with an NVIDIA TITAN Xp GPU and an Intel(R) Xeon(R) E5-2680 CPU. Our implementation of BOPO using PyTorch and trained models for each problem are available.111 https://github.com/L-Z-7/BOPO
@@
Exact Solver RL-based Improvement Traditional PDR RL SLL BOPO $B^{\prime}=128$ $B^{\prime}=512$
@@
5.1 Job-shop Scheduling Problem
Neural Model. Each JSP instance is represented as a disjunctive graph, a standard representation for scheduling problems. For details of disjunctive graphs, see Appendix A. We employ a neural model, named MGL, that combines a multi-layer graph attention network (GAT) (Veličković et al., 2018) encoder for computing node embeddings with a long short-term memory (LSTM) based (Hochreiter & Schmidhuber, 1997) context-attention decoder for predicting action probabilities using both embedding and context features. The complete architecture is detailed in Appendix B.
Training & Test. For evaluation, we use three standard JSP benchmarks: Taillard’s (TA) (Taillard, 1993), Lawrence’s (LA) (Lawrance, 1984), and Demirkol’s (DMU) (Demirkol et al., 1998). Each benchmark contains 8 different shapes with 10 instances per shape, except LA which has 5 instances per shape. We generate a training dataset of 30000 instances following SLIM (Corsini et al., 2024), consisting of 6 shapes $(n\times m)$ in $\{10\times 10,15\times 10,15\times 15,20\times 10,20\times 15,20\times 20\}$ with 5000 instances per shape. During training, we generate additional 100 different instances per shape from the same shape set for validation. We employ the Adam optimizer (Kingma & Ba, 2014) with learning rate $\eta=0.0002$ and train the neural model for 20 epochs. We set the solution number of hybrid rollout $B=256$, the number of filtered solutions $K=16$, batch size of $D=1$. During inference, we adopt both greedy rollout and sampling rollout with $B^{\prime}$ solutions.
[…]
@@ 5.2 Traveling Salesman Problem
Neural Model. BOPO adopts the same model as the typical POMO (Kwon et al., 2020), comprising an encoder with 6 Transformer layers and a decoder with a multi-head attention layer. To further demonstrate the universality of BOPO, it also employs the state-of-the-art INViT (Fang et al., 2024) model.
Training & Test. Following the NCO literature, we evaluate the proposed method on randomly generated instances with $n=20/50/100$ (denoted as TSP20/50/100), using models trained on corresponding problem shapes. Training instances are generated randomly from an uniform distribution. Additionally, we assess BOPO’s generalization capability on the out-of-distribution TSPLIB benchmark (Reinelt, 1991), and randomly generated instances with four different distributions, including uniform, cluster, explosion, and implosion distributions. We adopt the same hyperparameter configuration as POMO and INViT, respectively. During training, the number of filtered solutions is set to $K=8$. For TSP20 and TSP50, we set the hybrid rollout solution number $B=128$ and batch size $D=64$. For TSP100, we set $B=256,K=16$ and $D=48$ due to the memory limit. Specifically, we adopt BOPO to train INViT-2V, but due to memory constraints, we set $D=1,B=64$ (half of batch size in INViT), with $K=8$ and double epochs. For inference, we adopt greedy rollout with multiple start nodes and $\times$8 instance augmentation (denoted as aug.), consistent with POMO.
Baselines. We assess BOPO against both traditional methods and neural constructive methods. The traditional methods include exact solver Concorde and Gurobi, and LKH3, a powerful problem-specific heuristic. The neural constructive methods include POMO (Kwon et al., 2020), a widely-adopted RL-based backbone for advanced methods; DABL (Yao et al., 2024), a state-of-the-art SL-based method with data augmentation for routing problems; INViT (Fang et al., 2024), a state-of-the-art RL-based method; and SLIM (Corsini et al., 2024), a state-of-the-art SLL-based method applied to the POMO model.
Results on TSP. Results based on POMO are presented in Table 2 and Table 3. As shown in Table 2, BOPO outperforms other neural baselines (except DABL on TSP50) while delivering competitive solutions against traditional solvers, despite the latter requiring substantially more computational time. Compared with DABL which requires expensive labeled optimal solutions for SL, BOPO achieves comparable performance, demonstrating its superiority. Notably, our BOPO achieves superior performance over POMO on TSP100, even with fewer training epochs (700 vs. 2000). Generalization results on TSPLIB presented in Table 3 demonstrate BOPO’s significant advantages over other neural baselines when generalizing to out-of-distribution instances. Moreover, results based on INViT’s generalization to other distributions are shown in Figure 2, showing consistent superiority when implemented on different models.
@@
BOPO (aug.) 3.83 0.00 3.6s 5.69 0.01 6.6s 7.75 0.04 18.1s
Figure 2: Performance on TSP with different problem scales and distributions: (a) uniform, (b) cluster, (c) explosion, and (d) implosion. RL-original denotes the results reported in the INViT paper; RL-retrained denotes the results retrained with the same batch size as BOPO.
@@ 5.3 Flexible Job-shop Scheduling Problem
Neural Model. For an FJSP instance represented as a disjunctive graph, we adopt the MGL model (see Appendix B) to compute action probabilities, which is similar to the model for JSP.
Training & Test. For evaluation, we use FJSP instances from the LA benchmark (Lawrance, 1984). The benchmark includes e-data, r-data, and v-data, where each operation can be allocated to 1-2 machines, 1-3 machines, and 1-$m$ machines, respectively. Following Song et al. (2023), we generate 25000 FJSP instances with 5 different shapes: {$10\times 5\times 5$, $10\times 10\times 10$, $15\times 10\times 10$, $20\times 5\times 5$, $20\times 10\times 7$}. For each shape, 2500 instances belong to r-data and 2500 to v-data. We set $B=128$, $K=16$, and $D=1$ during training, and employ both greedy rollout and sampling rollout with $B^{\prime}$ solutions during inference.
Baselines. For FJSP, we compare our proposal with both representative traditional PDRs and state-of-the-art RL-based constructive methods: DNN (Yuan et al., 2024) using the actor-critic framework, HG (Song et al., 2023) utilizing heterogeneous graphs for instance representation, and RS (Ho et al., 2024) employing residual scheduling to remove finished operations.
Results on FJSP Benchmarks. As shown in Table 4, with greedy rollout, BOPO significantly outperforms most baselines, only marginally falling behind RS on LA e-data. When all methods use sampling rollout with 100 solutions, the proposed method even achieves the best performance across all cases. Notably, increasing the number of sampled solutions consistently reduces the optimality gap.
@@
Traditional PDR RL BOPO $B^{\prime}$=100 $B^{\prime}$=256 $B^{\prime}$=512
Benchmarks SPT MOR MWR DNN HG RS BOPO HG RS BOPO BOPO BOPO
@@
Shape Sampling Rollout Random Top-$K$ Bottom-$K$ Full Permutation Pairing BOPO
@@
5.4 Ablation Study
[…]
Figure 3: Training curves of different training paradigms for: (a) JSP and (b) TSP.
Effectiveness of the Best-anchored Preference Pair Construction Method. To validate our three-step preference pair construction method, we replace each step with a simpler alternative. We substitute sampling rollout for hybrid rollout, replace uniform filtering with random, top-$K$, or bottom-$K$ filtering, and use full permutation pairing instead of best-anchored pairing while maintaining the same total number of pairs. As shown in Table 5, BOPO’s performance deteriorates without (w/o) any of these components, demonstrating the necessity of our design. Notably, top-$K$ filtering leads to significantly worse performance, highlighting the importance of filtering worse solutions to create sufficient preference differences in pair construction. Additional analyses of hybrid rollout effectiveness are provided in Appendix E.
Superiority of the Objective-guided Preference Optimization Loss. To evaluate our proposed loss with a preference-based scaling factor, we compare it with other preference optimization losses: the classic DPO loss ($\mathcal{L}_{DPO}$) (Rafailov et al., 2023), the popular SimPO loss ($\mathcal{L}_{SimPO}$) (Meng et al., 2024), and a variant of BOPO loss without the scaling factor ($\mathcal{L}_{BOPO^{-}}$). Complete loss formulations are provided in Appendix C. For DPO, which requires a reference model to prevent excessive policy deviation, we use the old model from 10 episodes prior, similar to PPO, with the common hyperparameter setting $\beta=0.1$. For SimPO, we follow their standard parameters with $\beta=2$ and $\gamma=1$. As shown in Table 6, BOPO achieves the best performance across all benchmarks for both $B^{\prime}=128$ and $B^{\prime}=512$, with particularly significant improvements on the DMU benchmark, demonstrating the effectiveness of our loss design and its preference-based scaling factor. It is worth noting that SimPO performs poorly on DMU, where the test distribution diverges from training. This hints at potential overfitting caused by its target margin term $\gamma$.
@@ 5.4 Ablation Study
Benchmark $\mathcal{L}_{SimPO}$ $\mathcal{L}_{DPO}$ $\mathcal{L}_{BOPO^{-}}$ $\mathcal{L}_{BOPO}$ $\mathcal{L}_{SimPO}$ $\mathcal{L}_{DPO}$ $\mathcal{L}_{BOPO^{-}}$ $\mathcal{L}_{BOPO}$
@@ 5.5 Hyperparameter Study
BOPO has two crucial hyperparameters: the solution number of hybrid rollouts $B$ and the number of filtered solutions $K$. We analyze their individual effects by varying $B\in\{32,64,128,256,512\}$ and $K\in\{4,8,16,32\}$. Additionally, we explore their interaction by maintaining a fixed ratio $B/K=16$ while scaling both parameters $K\times B\in\{4\times 64,8\times 128,16\times 256,32\times 512\}$.
Effect of the Solution Number of Hybrid Rollouts. As shown in Figure 4(a), increasing the number of sampled solutions during training, i.e., larger $B$, improves solution quality by enhancing the probability of collecting higher-quality solutions. However, GPU memory consumption grows with $B$ due to parallel computation, raising costs. While the performance improves significantly up to $B=256$, further increases yield diminishing returns despite rising computational costs, making $B=256$ a reasonable choice. We also evaluate our method on TSP20/50 with $B=20/50$, the performance remains comparable to $B=128$ (detailed in Appendix F). This suggests that for small-scale problems, the optimal $B$ is low since the model can efficiently sample high-quality solutions.
@@ 6 Conclusion
In this work, we present BOPO, a preference optimization-based training paradigm for neural combinatorial optimization. By introducing best-anchored preference pair construction and a novel objective-guided pairwise loss function for COPs, our proposal achieves higher sample efficiency than mainstream RL and recent SLL paradigms. Extensive experiments on JSP, TSP, and FJSP demonstrate BOPO’s superior performance over state-of-the-art neural constructive methods, while requiring significantly less time to deliver solutions competitive with traditional problem-specific iterative heuristics. The proposed method requires neither expensive labels nor specialized design of the Markov Decision Process, making it easy to use in practice. More importantly, it establishes a general training paradigm that can be readily applied to various neural models for solving different COPs.
Although BOPO demonstrates superior sample efficiency compared with RL and SLL, one limitation is that it still requires a relatively large number of rollout solutions, similar to SLIM. This incurs moderate costs in collecting high-quality solutions for effective model learning. In future work, we will explore efficient ways to obtain high-quality solutions that facilitate model training. One promising direction is to leverage problem invariance and solution symmetry to efficiently generate diverse high-quality training data. Another direction is to efficiently enhance solution quality by incorporating problem-specific heuristics during training, providing better learning signals for the model.
Acknowledgements
This work is supported by the National Natural Science Foundation of China (62472461) and the Guangdong Basic and Applied Basic Research Foundation (2024A1515010871, 2025A1515010129).
@@ A.1 Job-shop Scheduling Problems
Disjunctive Graph. JSP can be represented using a disjunctive graph $G=(V,A,E)$. In this graph, the node set $V=\{O_{j}\;|\;O_{j}\in\mathcal{O}\}$ represents operations, the directed edge set $A=\{O_{l_{ij}}\rightarrow O_{l_{i(j+1)}}\;|\;O_{l_{ij}},O_{l_{i(j+1)}}\in\mathcal{O}\}$ indicates precedence constraints between successive operations ($O_{l_{ij}}\rightarrow O_{l_{i(j+1)}}$), and the disjunctive (undirected) edge set $E$ connects operations performed on the same machine. A feasible solution is obtained by assigning directions to the undirected edges in $E=\{O_{j}\leftrightarrow O_{j^{\prime}}\;|\;O_{j},O_{j^{\prime}}\in\mathcal{O},M_{j}=M_{j^{\prime}}\}$, resulting in a directed acyclic graph.
Features of JSP. Following previous works (Zhang et al., 2020; Corsini et al., 2024), we define two types of features. The first type consists of static state features $\bm{s}_{j}$ assigned to each operation $O_{j}$ in the node set $V$, while the second type comprises contextual features $\bm{c}_{i}$ based on job $J_{i}$ information at step $t$. The state features $\bm{s}_{j}$ are fed into the encoder to compute embeddings for each node (operation), while the contextual features $\bm{c}_{i}$ are combined with these embeddings to jointly contribute to the prediction. Table 7 details the state features $\bm{s}_{j}\in\mathbb{R}^{15}$, and Table 8 details the contextual features $\bm{c}_{i}\in\mathbb{R}^{11}$.